Zinc metal reacts with hydrochloric acid according to the following balanced equation.
Zn(s) + 2HCl(aq) -> ZnCl2(aq) + H2(g)

When 0.104g of Zn(s) is combined with enough HCl to make 54.5 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 22.4 Celcius to 24.7Celcius.

Find the delta Hrxn (in kJ/mol) for this reaction as written. (Use 1.0g/mL for the density of the solution and 4.18 J/g *Celcius as the specific heat capacity.)
Answer #1

Mass of solution = volume x density

= 54.5 x 1.0 = 54.5 g

Heat released by reaction = heat absorbed by solution

= mass x specific heat x temperature change of solution

= 54.5 x 4.18 x (24.7 - 22.4)

= 523.963 J

Moles of Zn = mass/molar mass of Zn

= 0.104/65.41 = 0.001590 mol

ΔH = -heat released by reaction/moles of Zn

= -523.963/0.001590

= -3.30 x 105 J/mol = -330 kJ/mol

Note that ΔH is negative as reaction is exothermic and heat is released